A car traveling at 91.0 km/h approaches the turnoff for a restaurant 30.0 m ahead. If the driver slams on the brakes with an acceleration of -6.40 m/s^2, what will be her stopping distance?


1
Expert's answer
2021-11-04T10:17:07-0400

We can find the stopping distance of the car from the kinematic equation:


"v^2=v_0^2+2ad,""0=v_0^2+2ad,""d=\dfrac{-v_0^2}{2a}=\dfrac{-(25.28\ \dfrac{m}{s})^2}{2\times(-6.4\ \dfrac{m}{s^2})}=50\ m."